is there any shortcuts ti remember the all trigonometry formulas

Grade:11

6 Answers

Rathod Shankar AskiitiansExpert-IITB
13 years ago

There is Always a time in Exam for your Questions.

1) Yes there is Always a way to solve trignometry questions by shorcut. Just substitute any angle in 2 sides if 2 angles satisfies both the equations then try for another angle  by repeating this u will know the answers.

2) but it is very good if u dont use shortcuts. if u do that after crack JEE u can do nothing and u have to hace concept. so build your concepts

 

Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.
Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar..

swapnil sudhir
13 years ago

Dear Nitin ,

To learn simple Trigonometric formulas  the most commonly used technique is to revise the questions many times so as to revise the formulae. Learn The formulaes Early in the morning as they help to cross out your problems. Remember  trigonometry scores about 30% in iit jee.

SAGAR SINGH - IIT DELHI
13 years ago

Dear nithin,

Basic trigonometric ratios will be defined using the right angled triangle ABC given below , to later on move to trigonometry formulas.

Trigonometry Formulas

Trigonometric Ratios

Sin X = Opposite Side/ Hypotenuse

Cos X = Adjacent Side/ Hypotenuse

Sec X = 1/Cos X ; Csc X = 1/ Sin X; Tan X = Sin X/ Cos X; Cot X = Cos X/Sinx = 1/ Tan X

Trigonometric Formulas Sub topics

Pythagorean identities

Sin2X + Cos2 X = 1

Tan 2X + 1 = Sec2X

Cot2X + 1 = Csc2X

 

We are all IITians and here to help you in your IIT JEE preparation.

All the best.

 If you like this answer please approve it....

win exciting gifts by answering the questions on Discussion Forum

 

Sagar Singh

B.Tech IIT Delhi

sagarsingh24.iitd@gmail.com

ayush kumar
10 years ago

hi bhaiya i m also studying for iit , and wanted to know that are there any tricks to solve trignometry option wise. and could u plz give me some tips and experiences. 

looking forward for your reply.

plzz ans me at my email id ayushshivrain@gmail.com

wfwf
8 years ago
fuck yourselves
 
wfwf
8 years ago
any